A friend is planning to purchase a new car for $35,000 and intends to borrow money to finance the entire purchase. She has asked you to compare the following options and to let her know which one you think is best:

option I: $5000 price reduction, plus a four-year loan, with monthly payments. The annual interest rate on the loan will be 5%.

Option II: A zero interest loan, with a 4-year term and monthly payments.

Use formulas and show step by step with calculations
Answer #1

We need to compare the total payments under the two options

Option 1:

Using financial calculator
Input:

PV = -30000

N= 4*12 = 48

I/Y= 5/12

Solve for PMT as 690.88

Total payments = 690.88*48 = $ 33162.18

Option 2

Since there is no interest, Total payments = amount of loan = 35000

Option 1 is better since the total payments are lesser.